In a chemistry class, 12L of a 12% alcohol solution must be mixed with a 20% solution to get a 14% solution. How many liters of the 20% solution are needed?
How do I set the problem up and solve?
I'm so confused and would appreciate any help. Thanks so much.

Answer by rwm(914)

.12*12+.2x=.14*(12+x)
.12*12+.20x=.14*12+.14x
.06x=.02*12
6x=2*12
x=4